OpenStudy (bananas1234):

what's h(-3) given h(x) = -/2 - x/?

OpenStudy (blazeryder):

PLug in -3 where the x is. And then solve.

OpenStudy (bananas1234):

\[h(x) = -|2 - x|\]

OpenStudy (tkhunny):

Substitution. That's how the notation works. If f(x) = 3 + x, then f(2) = 3 + 2 If h(x) = x - 9/x, then h(4) = 4 - 9/4

OpenStudy (bananas1234):

so 1? still kinda confused

OpenStudy (tkhunny):

Did you substitute '-3' everywhere is says 'x' in the definition?

OpenStudy (bananas1234):

-5?

OpenStudy (tkhunny):

Are you guessing? Please show your work.

OpenStudy (bananas1234):

no im not guessing. -/2 - (-3)/ 2- (-3) = 5 -5

OpenStudy (tkhunny):

Excellent. Go with confidence.

OpenStudy (bananas1234):

thank you